OpenStudy (anonymous):

Simplify and determine the domain. ?? x2 + 5x +6 x2 - 3x -10

sam (.sam.):

All real

OpenStudy (anonymous):

i am going to make a guess that this is supposed to be \[\frac{x^2 + 5x +6}{x^2 - 3x -10}\]

OpenStudy (anonymous):

in which case you can factor and cancel perhaps, \[\frac{x^2 + 5x +6}{x^2 - 3x -10}=\frac{(x+2)(x+3)}{(x-5)(x+2)}\] so cancel but with the stipulation that \(x\neq 5,x\neq -2\)